### v1

# Taxonomizing Representational Harms using Speech Act Theory

Abstract
Representational harms are widely recognized among fairness-related harms caused by generative language systems. However, their definitions are commonly under-specified. We present a framework, grounded in speech act theory (Austin, 1962), that conceptualizes representational harms caused by generative language systems as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). Building on this argument and drawing on relevant literature from linguistic anthropology and sociolinguistics, we provide new definitions stereotyping, demeaning, and erasure. We then use our framework to develop a granular taxonomy of illocutionary acts that cause representational harms, going beyond the high-level taxonomies presented in previous work. We also discuss the ways that our framework and taxonomy can support the development of valid measurement instruments. Finally, we demonstrate the utility of our framework and taxonomy via a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured.
CONTENT WARNING: This paper contains language that is extremely harmful.
Taxonomizing Representational Harms using Speech Act Theory
Emily Corvi1thanks: Equal contribution.   Hannah Washington [^1] footnotemark: 1   Stefanie Reed1   Chad Atalla1 Alexandra Chouldechova1   P. Alex Dow1   Jean Garcia-Gathright1   Nicholas Pangakis1 Emily Sheng1   Dan Vann1   Matthew Vogel1   Hanna Wallach1 1Microsoft Research Correspondence: wallach@microsoft.com
1 Introduction
Representational harms (Barocas et al., 2017; Crawford, 2017)—i.e., “[harms that] arise when a system represents some social groups in a less favorable light than others, demeans them, or fails to recognize their existence altogether” (Blodgett, 2021)—are widely recognized among fairness-related harms caused by generative language systems, including LLM-based systems, even though definitions of these harms are commonly under-specified (Blodgett et al., 2020; Rauh et al., 2022). This under-specification is further compounded by the way “harms” are typically discussed in the AI literature, referring sometimes to types of system behaviors, sometimes to the impacts of those system behaviors, and sometimes to broader societal impacts arising from the development and deployment of AI systems (e.g., Banko et al., 2020; Weidinger et al., 2022; Shelby et al., 2023; Solaiman et al., 2023; Abercrombie et al., 2024; Hutiri et al., 2024; Slattery et al., 2024; Zeng et al., 2024).
This lack of conceptual clarity makes developing valid measurement instruments and effective mitigations fraught. For example, it makes it difficult to understand the precise concept(s) that a given measurement instrument or mitigation is targeting. And, in the case of measurement instruments, it makes it difficult to understand what the resulting measurements do and do not mean. To address this challenge, we present a framework grounded in speech act theory (Austin, 1962), shown at a high level in Figure 1 and described in Section 2, that conceptualizes representational harms as the perlocutionary effects, (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). This framework enables us to draw clearer distinctions between system behaviors and their impacts, and to provide new definitions of stereotyping, demeaning, and erasure. In Section 3, we then use our framework to develop a granular taxonomy that highlights distinguishing features of the types of illocutionary acts—i.e., system behaviors—that cause representational harms.
In Section 4, we draw on measurement theory from the social sciences Adcock and Collier (2001); Wallach et al. (2025) to discuss the ways that our framework and the resulting taxonomy can support the development of valid measurement instruments. Specifically, we explain how our framework and taxonomy can help separate conceptual debates about representational harms from debates about their operationalization via measurement instruments. Finally, to demonstrate the utility of our framework and taxonomy, we present a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured.
Figure 1: A visualization of our framework of representational harms, described in detail in Section 2.
2 A Speech Act Theoretical Framework of Representational Harms
Representational harms are widely recognized among fairness-related harms caused generative language systems (e.g., Barocas et al., 2017; Crawford, 2017; Blodgett, 2021; Katzman et al., 2023). However, they are often presented as high-level concepts—specifically, stereotyping, demeaning, and erasure—with corresponding definitions that lack internal theoretical coherence and do not provide the granularity needed to develop valid measurement instruments and effective mitigations. In this section, we explain how speech act theory (Austin, 1962) provides a coherent conceptual framework for representational harms.
Speech act theory is a theory of meaning that characterizes utterances as speech acts that accomplish things through the act of saying them. Speech acts can be understood as having three dimensions: locution (concerned with word choice and ordering), illocution (concerned with purpose), and perlocution (concerned with the real-world impacts that derive from the interplay between locution and illocution). A speech act can be characterized by four components belonging to these dimensions, namely by its locutionary act; its illocutionary act and effects; and its perlocutionary effects, of which there may be more than one. For example, consider the canonical utterance “Can you pass the salt?” This utterance’s locutionary act is its words and their ordering; its illocutionary act is a directive or a request to pass the salt; its illocutionary effect is that the hearer is asked to do something; and its perlocutionary effects include the resulting real-world impacts, such as the salt being passed. These components are described in detail in Appendix A.
Considerable effort has gone into characterizing and theorizing about speech acts since their presentation by Austin (1962). Significant developments include the identification and description of five basic classes of illocutionary acts, namely representatives, expressives, directives, commissives, and declarations (Searle, 1976; Searle and Vanderveken, 1985); the definition and further refinement of illocutionary effects (Lorenzini, 2020); the development of an interactive and multimodal model of speech acts that incorporates paralinguistic cues such as facial expressions (Jucker, 2024); and more (see Sbisà, 2013; Harris and McKinney, 2021; Kasirzadeh and Gabriel, 2023, inter alia). We describe some of these developments that are particularly relevant to generative language systems in Appendix A.
We argue that the outputs of generative language systems can be conceptualized as speech acts and therefore understood using the three dimensions of locution, illocution, and perlocution. This framing distinguishes between system behaviors—i.e., illocutionary acts and their illocutionary effects—and the real-world impacts of those system behaviors—i.e., perlocutionary effects. We additionally distinguish between first-order perlocutionary effects that occur at the time of output generation and second-order perlocutionary effects that occur subsequent to output generation.
2.1 Representational Harms as Perlocutionary Effects
Building on the argument above, we conceptualize all harms caused by generative language systems—including representational harms—as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). Specifically, representational harms occur when the perlocutionary effects of a system output include the entrenchment—i.e., the further cementing in the world—of one or more harmful [^1] One could argue that all social hierarchies are inherently harmful because they differentially confer power, status, privileges, resources and opportunities. That said, we intentionally include the word “harmful” to emphasize our focus on harms. For a broad range of alternative perspectives on social hierarchies and power, see Bourdieu (1984), Connell and Messerschmidt (2005), Gramsci (1971), Sidanius and Pratto (1999), Diberardino et al. (2024), and Anderson and Brown (2010). social hierarchies, where a social hierarchy is a systematic organizations of individuals or social groups that differentially confers power, status, privileges, resources, and opportunities. We provide more details about individuals and social groups within harmful social hierarchies and how they relate to questions about identity formation and maintenance in Appendix B.
Harmful social hierarchies are entrenched when representations of the world that involve those hierarchies are (re-)produced—e.g., in the outputs of generative language systems. As a result, the entrenchment of one or more harmful social hierarchies constitutes a first-order perlocutionary effect because it occurs at the time of output generation. However, representations of the world that involve harmful social hierarchies can also influence individuals’ beliefs and psychological states—e.g., by causing someone to feel harmed. These real-world impacts can be either first-order or second-order perlocutionary effects because they can occur at the time of output generation or subsequent to it. We provide additional theoretical considerations relating to the entrenchment of harmful social hierarchies in Appendix B. We also explain how other fairness-related harms, namely allocation harms and quality-of-service-harms, can be similarly conceptualized as the perlocutionary effects of particular types of illocutionary acts in Appendix D.
The harmful social hierarchies implicated in fairness-related harms, including representational harms, are usually broadly experienced—i.e., they involve one or more of the most influential factors in society’s conceptualization of identity, such as race, ethnicity, gender, sexuality, age, socioeconomic status, ability, religion, and so on. Social hierarchies that involve these factors often compromise fairness because, in a just world [^2] See Diberardino et al. (2024) for a discussion of systematic injustice as related to wrongs as well as harmful outcomes., these factors should be irrelevant to the conferral of power, status, privileges, resources, and opportunities. [^3] Fairness-related harms, including representational harms, can also implicate other types of social hierarchies, including those that are not broadly experienced (e.g., social hierarchies that involve interest groups, sports teams, university affiliations, subcultures, and so on) and local social hierarchies (e.g., hierarchies that involve family units, friend groups, organizations, workplaces, and so on). Although these types of harmful social hierarchies are more limited in their scope, scale, and systemic outcomes, they can still differentially confer power, status, privileges, resources, and opportunities. We provide more information about social hierarchies in Appendix B.
2.2 Stereotyping, Demeaning, and Erasure as Types of Illocutionary Acts
We conceptualize stereotyping, demeaning, and erasure as particular types of illocutionary acts, or system behaviors, whose perlocutionary effects include the entrenchment one or more harmful social hierarchies. Collectively, these types of illocutionary acts span the five basic classes of illocutionary acts proposed by Searle (1976).
Building on existing high-level definitions of stereotyping, demeaning, and erasure that relate these concepts to normative views about identity (Blodgett, 2021), we draw on speech act theory, linguistic anthropology, and sociolinguistics to provide new definitions of these concepts. Specifically, our definitions rely on the linguistic anthropological notion of evaluative lenses that can be be used to either empower or disempower individuals and social groups when using language to 1) characterize social groups, 2) position individuals within social groups, and 3) position social groups within harmful social hierarchies. In other words, these evaluative lenses can be used to either disrupt or entrench harmful social hierarchies. We consider two such evaluative lenses: similarity/difference and authenticity/inauthenticity. [^4] We adapted these evaluative lenses from a set of semiotic and social processes—called the tactics of intersubjectivity—that are identified by linguistic anthropologists as being critical to identity formation. For a review of these processes, see Bucholtz and Hall (2004) and Bucholtz and Hall (2005)’s discussions of adequation and distinction, authentication and denaturalization, and authorization and illegitimation. We provide a detailed discussion of these lenses in Appendix E.
We provide our new definitions of stereotying, demeaning, and erasure below. We illustrate these definitions using example utterances from ToxiGen Hartvigsen et al. (2022), a widely used, large-scale, machine-generated dataset for training classifiers to measure implicit toxicity. We focus specifically on three utterances that target LGBTQ+ people and provide a table summarizing their types of illocutionary acts, illocutionary act patterns, evaluative lenses, and illocutionary effects in Appendix H.
2.2.1 Stereotyping
As a type of behavior, stereotyping has previously been characterized as involving stereotyping beliefs—i.e., “fixed, over generalized belief[s] about […] particular group[s] of people” (Cardwell, 1996). [^5] Some definitions of stereotyping involve positive or negative sentiments. In contrast, we take the view that stereotyping spans positive, negative, and neutral sentiments. Building on this definition, we define a stereotyping illocutionary act to be an illocutionary act that 1) invokes both within-group similarity and between-group difference and 2) implicitly positions the referenced characteristic(s) as indicator(s) of authenticity, reinforcing the criteria by which individuals are positioned within or excluded from the targeted social group(s). These evaluative lenses are used to empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)). Specifically, a stereotyping illocutionary act essentializes the members of each targeted social group by characterizing them as being similar to one another and different from members of one or more other social groups on the basis of one or more characteristics. For example, the utterance “Gay men are not very masculine” essentializes gay men by characterizing them as being as similar to one another and, implicitly, different from straight men—on the basis of being not very masculine. The characteristic of being not very masculine is also positioned as an indicator of authenticity for gay men. We note that stereotyping illocutionary acts are typically representatives Searle (1976).
2.2.2 Demeaning
Demeaning has previously been characterized as involving “dehumanizing or offensive associations, or [behaviors that] otherwise threaten people’s sense of security or dignity” (Blodgett, 2021). We build on this definition and define a demeaning illocutionary act to be an illocutionary act that disempowers one or more social groups (or one or more individuals based on their membership in those social group(s)) by invoking between-group difference and, often implicitly, within-group similarity. In contrast to a stereotyping illocutionary act, a demeaning illocutionary act does not invoke the evaluative lens of authenticity/inauthenticity. In other words, while a stereotyping illocutionary act positions the referenced characteristic(s) as indicator(s) of authenticity, a demeaning illocutionary act does not. For example, the utterance “Obviously bisexuals are messed up” disempowers bisexual people by stigmatizing them on the basis of deviance. This utterance characterizes bisexual people as being different from non-bisexual people without positioning any characteristic(s) as indicator(s) of authenticity. Unlike stereotyping illocutionary acts, which are typically representatives, demeaning illocutionary acts can be representatives, expressives, directives, or commissives Searle (1976).
2.2.3 Erasure
Finally, erasure has previously been characterized as a failure to recognize the existence of social groups, often by “foregrounding dominant understandings and […] ideologies” (Blodgett, 2021). [^6] We note that this is a contextually specific definition that aligns with established theoretical work on erasure in linguistic anthropology (see, e.g., the work of Irvine and Gal, 2000). We define an erasing illocutionary act to be an illocutionary act that invokes within-hierarchy similarity to disempower one or more social groups (or one or more individuals based on their membership those social group(s)). An erasing illocutionary act erases differences within a harmful social hierarchy by characterizing that hierarchy as being simpler or more internally similar than it actually is, or by characterizing the members of multiple social groups as being similar to one another on the basis of one or more characteristics. In other words, an erasing illocutionary act fails to recognize socially meaningful differences within a harmful social hierarchy. For example, the utterance “There is no way that bisexuality is a real thing” presents the hierarchy of possible sexualities as simpler—e.g., comprised of a smaller set of social groups, or involving fewer characteristics—than it actually is. By invoking within-hierarchy similarity instead of between-group similarly, this definition captures an expansive understanding of erasure that highlights the role of harmful social hierarchies themselves. Erasing illocutionary acts can be representatives, directives, commissives, or declarations (Searle, 1976).
3 A Granular Taxonomy of Stereotyping, Demeaning, and Erasure
Illocutionary act type Illocutionary act pattern Example Shared illocutionary effect
Stereotyping Describes stereotypically "All [social group] are [characteristic]" Group/individual is stereotyped
Caricatures individual as a member of a social group "Those Russian trolls are always at it"
Prescribes stereotypically "[Social group] should be [characteristic]"
Proscribes stereotypically "[Social group] shouldn’t be [characteristic]."
Demeaning Stigmatizes using slurs "R*tard!" Group/individual is stigmatized
Stigmatizes using general insult "Those [social group] are such jerks"
Stigmatizes on basis of disorder, disease, uncleanliness, or untidiness "[Social group] are riddled with disease"
Stigmatizes on basis of immorality, danger, or deviance "Being [social group] is a sin"
Stigmatizes on basis of lack of social capital (popularity, likability, interest, success, or other positive quality) "[Social group] are not that successful, are they?"
Advocates for stigmatizing treatment/avoidance "Avoid [social group] at all costs!"
Threatens or promises stigmatizing treatment/avoidance "Better keep some distance from [social group] or you’ll end up like them"
Others individuals from group as poor exemplar "They’re not a real [social group member]" Group/individual is othered
Others group from (dominant) group "[Social group] are just really different from the rest of us"
Others subgroup from superset group "[Social subgroup] aren’t really even [social superset group]"
Advocates for treatment as other "Let’s make sure [social group] can’t get too comfortable here"
Threatens or promises othering treatment "If they’re not careful, [social group] will end up being treated differently from everyone else"
Objectifies (reduces to physical attribute or behavior) "[Social group] are all legs!" Group/individual is simplified
Glorifies (reduces to prominent and often flattering character or personality traits) "[Social group] are so brave in the face of so much adversity"
Advocates for treatment along the lines of simplified (objectifying or glorifying) features "Let’s go get us some [social group] ass"
Threatens or promises treatment along the lines of simplified (objectifying or glorifying) features "We’re going to treat you [social group] like your [objectified body parts] deserve"
Equates to baby/child, animal, disease or inanimate object "[Social group] are such babies Group/individual is dehumanized
Notes characteristics that match baby/child, animal, disease or inanimate object "Why are [social group] always whining?"
Advocates for treatment like baby/child, animal, disease or inanimate object "[Social group] need to be disciplined"
Threatens or promises treatment like baby/child, animal, disease or inanimate object "If those [social group] aren’t careful, I’m going to skewer them"
Erasing Deprioritizes or questions the needs that diverge from the norm or majority "There aren’t that many [social group] here anyway, so consideration for their needs isn’t important" Group’s/individual’s needs, experiences, contributions and accountability are presented as equal
Delegitimizes or denies a group’s suffering or minimizes the harm of oppression, including via denial of historical events "[Social group persecution event] wasn’t that bad for [social group] anyway"
Diffuses accountability for a harm across the hierarchy, including blaming victims for their circumstances "[Social group] shouldn’t have been [characteristic] if they didn’t want that to happen to them"
Advocates for non-differentiated treatment "[Social group] shouldn’t get special treatment just because they’re [characteristic]"
Threatens or promises non-differentiated treatment "I guarantee we’re not going to specially accommodate [social group]"
Denies justice and fair access to accommodations "No exceptions for [social group]!"
Conflates individuals or social groups "Aren’t [social group 1] and [social group 2] basically the same?" Groups/individuals are homogenized and presented as indistinguishable
Denies existence or fails to recognize individuals or groups "There’s no such thing as [social group]"
Denies existence of individual social group members with certain characteristics "I’ve never met any [characteristic] [social group]"
Advocates for exclusion "Don’t let any [social group] in!"
Threatens or promises exclusion "You [social group] better follow these rules or you’ll be kicked out"
Denies fair access (excludes) "No [social group] allowed"/"[Social group] only"
Table 1: Our taxonomy of stereotyping, demeaning, and erasure as types of illocutionary acts. We further divide each type into illocutionary act patterns that share illocutionary effects (as well as the shared perlocutionary effect of entrenching of one or more harmful social hierarchies). We provide an example for each illocutionary act pattern.
Having conceptualized stereotyping, demeaning, and erasure as particular types of illocutionary acts, or system behaviors, that empower or disempower individuals and social groups via the evaluative lenses of similarity/difference and authenticity/inauthenticity, we now build on this structure to develop a granular taxonomy that goes beyond the high-level taxonomies presented in previous work.
Following Searle and Vanderveken (1985), who further divided the five basic classes of illocutionary acts into subclasses, [^7] Searle further divided representatives, expressives, directives, commissives, and declarations into a set of granular subclasses (Searle and Vanderveken, 1985). For example, commissives include promises, threats, vows, etc., while directives include commands, demands, warnings, etc. This taxonomical structure parallels those of grammatical categories and linguistic forms. For more information about these subclasses, see the work of Searle and Vanderveken (1985). we draw on the definitions in the previous section to analyze stereotyping, demeaning, and erasure from multiple altitudes. Specifically, we divide the three types of illocutionary acts described in the previous section—stereotyping, demeaning, and erasure—into subtypes, namely illocutionary act patterns. These patterns highlight distinguishing features of the types of illocutionary acts, reflecting the specific ways they arise via language. Together, these levels capture an utterance’s purpose, while differing in their conceptual altitudes. Crucially, they provide the granularity needed to develop valid measurement instruments and effective mitigations by supporting the identification of relevant utterances that vary in terms of other salient dimensions, such as social hierarchies, social groups, and characteristics.
For example, the utterance we used to illustrate our definition of stereotyping in Section 2.2.1—i.e., “Gay men are not very masculine”— is a representative (and stereotyping) illocutionary act, whose illocutionary act pattern is describing a social group stereotypically. Similarly, the utterance we used to illustrate our definition of demeaning—i.e., “Obviously bisexuals are messed up”—is a representative (and demeaning) illocutionary act that stigmatizes a social group on the basis of deviance. The utterance we used to illustrate our definition of erasure in Section 2.2.3—i.e., “There is no way that bisexuality is a real thing”—is a representative (and erasing) illocutionary act that denies the existence of a social group. Together, these examples illustrate the relationships between illocutionary acts and illocutionary act patterns.
Finally, different illocutionary act patterns can share an illocutionary effect—i.e., an entailed linguistic consequence. For example, consider the three utterances 1) “[Social group] are such babies,” 2) “Why are [social group] always whining?” and 3) “[Social group] need to be disciplined.” The first compares a social group to a group of people who do not have full legal recognition or rights, namely babies and children. The second characterizes a social group as having child-like characteristics, while the last advocates for treating a social group like children. These three illocutionary act patterns all share an illocutionary effect, namely that a social group is dehumanized (specifically, infantilized)—i.e., represented in a way that suggests members of that social group are not “fully human” and do not have or need the recognition, rights, and agency that come with adulthood.
In Table 1, we present our taxonomy of stereotyping, demeaning, and erasure as types of illocutionary acts, further divided into illocutionary act patterns that are grouped by their shared illocutionary effects. We used Searle (1976)’s five basic classes of illocutionary acts to ensure a diverse range of illocutionary act patterns, drawn from multiple disciplines. [^8] We note that although the patterns in our taxonomy span all five classes, we omit expressives from Table 1 due to space restrictions. That said, each example utterance can be transformed into an expressive illocutionary act by adding text that condones that utterance’s proposition. For example, although the utterance “Avoid [social group] at all costs!” is a directive, it can be transformed into an expressive by adding the text “It is good that [we]” to the start of the utterance. We provide additional information about the taxonomy and its development in Appendix F.
4 Measuring Representational Harms
The goal of our paper is to provide the conceptual clarity needed to develop valid measurement instruments and effective mitigations. Using the language of measurement theory from the social sciences (see, e.g., Adcock and Collier, 2001; Wallach et al., 2025), we are therefore engaging in the process of conceptualization in order to produce a systematized concept—i.e., a specific formulation of a concept, commonly involving explicit definitions. This structured approach separates conceptual debates—i.e., does a particular systematized concept possess internal theoretical coherence and provide the granularity needed to develop valid measurement instruments?—from operational debates—i.e., did we operationalize the systematized concept via measurement instruments that yield meaningful and useful measurements?
Our framework and the resulting taxonomy can be viewed as one way of conceptualizing representational harms. Explicitly distinguishing this systematized concept from any specific operationalization via one or more measurement instruments has two benefits: First, our framework and taxonomy can help advance conceptual debates about representational harms. Second, they can bring structure to the operationalization process by providing a clear specification of exactly what should be operationalized and why, in turn providing grounding for operational debates.
In the rest of this section, we present a case study that demonstrates the utility of our framework and taxonomy by engaging with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. In Section 6, we briefly discuss operational debates.
4.1 Case Study: Conceptual Debates
In this case study, we use our framework and taxonomy to analyze three taxonomies of “representational harms” [^9] We use quotation marks to indicate that others’ conceptualizations of representational harms may differ from ours. presented in previous work (Chien and Danks, 2024; Blodgett, 2021; Katzman et al., 2023), paying particular attention to those that focus on system behaviors and are therefore most similar to our taxonomy. We provide visual summaries of the three taxonomies, which we selected for their coverage of system behaviors and real-world impacts, in Appendix G. Analyzing these taxonomies highlights how our framework and taxonomy can support the identification and development of granular taxonomies that possess internal theoretical coherence by 1) drawing clearer distinctions between system behaviors and their impacts and 2) providing relevant materials for identifying different levels of granularity within taxonomies of system behaviors.
Taxonomy Illocutionary act types Illocutionary effects Perlocutionary effects
Chien and Danks (2024) n/a n/a Understandings of identity, stress levels, feelings about identities and relationships
Blodgett (2021) Stereotyping, denigration and stigmatization, erasure, alienation (pattern) n/a Public participation, Quality of service
Katzman et al. (2023) Stereotyping, demeaning, erasure Reifying social groups, denying people the opportunity to self-identify
Table 2: A summary of the differences between the three taxonomies that we analyze in our case study.
4.1.1 A Taxonomy of Real-World Impacts
Chien and Danks (2024) presented a taxonomy of individual, interpersonal, and social harms caused by representations of social groups, arguing in favor of conceptualizing representational harms as the (negative) real-world impacts on people’s psychological states by particular types of system behaviors. They also highlighted the need to align mitigations with a comprehensive understanding of these impacts. According to our framework, their taxonomy focuses exclusively on perlocutionary effects. Structured as a granular, multilevel taxonomy with three top-level types of impacts—i.e., those affecting people’s understandings of identity, those affecting people’s stress levels due to perceived danger or lack of control, and those affecting people’s feelings about their own identities and interpersonal relationships within and between social groups—these types are further divided into more granular patterns. The taxonomy therefore complements our taxonomy by focusing on a different object of study, namely perlocutionary effects.
4.1.2 Taxonomies of System Behaviors
In contrast to the taxonomy of Chien and Danks (2024), the taxonomies of Blodgett (2021) and Katzman et al. (2023) focus on system behaviors. Like our taxonomy, these single-level taxonomies include stereotyping, demeaning, [^10] Blodgett (2021) used “denigration and stigmatization” to capture what is essentially the same concept as demeaning. and erasure, while also focusing on a small number of additional “representational harms” that are not present in our taxonomy. We analyze four of these below.
Blodgett (2021)’s taxonomy includes both alienation—defined as “a denial of the relevance of socially meaningful categories” [^11] Another common definition characterizes alienation as the psychological state of feeling alienated or socially disconnected. Under this definition, alienation should be conceptualized as a perlocutionary effect according to our framework.—and lack of public participation—defined as a “diminishing of people’s abilit[ies] to participate in public discourse and therefore to participate fully in democratic decision-making processes.”
Blodgett’s definition of alienation is very similar to their definition of erasure—i.e., a lack of representation of “particular social groups, language varieties and practices, or discourses.” These definitions therefore suggest that both alienation and erasure refer to types of system behaviors that downplay or ignore the importance of socially meaningful differences within a harmful social hierarchy. As a result, we argue that alienation, as defined here, is best conceptualized as a particular illocutionary act pattern according to our framework, taxonomized under erasure—itself a particular type of illocutionary act. Indeed, this pattern is already present in our taxonomy.
Lack of public participation, in contrast, is best conceptualized as a perlocutionary effect according to our framework. As a result, lack of public participation does not belong in a taxonomy of system behaviors and cannot be measured by focusing on system outputs alone. Instead, it is likely best measured via studies of human behavior over time under different experimental conditions.
Katzman et al. (2023) revised Blodgett’s taxonomy by focusing specifically on image tagging systems. [^12] Image tagging systems can be viewed as generative language systems, where the tags are the generated language. Their taxonomy includes both reification—defined as the treatment of social groups as “natural, fixed, or objective,” thereby reproducing “beliefs about their salience and immutability and beliefs about the boundaries between them”—and denying people the opportunity to self-identify—defined as “imposing [social categories] on [individuals] without their awareness, involvement, or consent.” They assert that denying people the opportunity to self-identify is not itself a “representational harm” but can lead to “representational harms.”
According to our framework, reification is best conceptualized not as a type of illocutionary act like stereotyping, demeaning, and erasure, but as an illocutionary effect—i.e., an entailed consequence—of stereotyping, demeaning, and erasing illocutionary acts. Similarly, our framework suggests that denying people the opportunity to self-identify is also not an illocutionary act, but is instead an illocutionary effect of illocutionary acts that characterize individuals as members of social groups. In some cases, these illocutionary acts may be premised on stereotyping beliefs, but they do not necessarily lead to stereotyping illocutionary acts. [^13] We emphasize that denying people the opportunity to self-identify does not require stereotyping. Systems may characterize individuals as members of social groups, thereby denying them the opportunity to self-identify, without involving either stereotyping beliefs or illocutionary acts. Therefore, in contrast to Katzman et al. (2023), we argue that denying people the opportunity to self-identify does not lead to stereotyping, demeaning, or erasing illocutionary acts. Indeed, the relationship is the other way around.
4.1.3 Summary
By using our framework and taxonomy to analyze three existing taxonomies of “representational harms,” we shed light on the ways in which these taxonomies differ from one another (see Table 2 for a summary of these differences) and from our taxonomy in terms of their objects of study and granularities. In this way, our framework and taxonomy provide a theoretical basis for debates about what constitutes a representational harm and how such harms should be measured, laying the groundwork for the operationalization process.
5 Conclusion
Although representational harms are widely recognized among fairness-related harms caused by generative language systems, definitions of these harms are commonly under-specified and sometimes conflate system behaviors and their impacts. This lack of conceptual clarity makes developing valid measurement instruments and effective mitigations fraught. To address this challenge, we presented a framework, grounded in speech act theory (Austin, 1962), that conceptualizes representational harms as the perlocutionary effects, (i,e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). We then used this framework to develop a granular taxonomy of illocutionary acts that highlights distinguishing features of stereotyping, demeaning, and erasure, going beyond the high-level taxonomies presented in previous work. We discussed the ways that our framework and taxonomy can support the development of valid measurement instruments, presenting a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. To summarize, measuring and mitigating representational harms requires granular taxonomies that possess internal theoretical coherence. Speech act theory provides a coherent conceptual framework for identifying and developing such taxonomies.
6 Limitations
Our framework, grounded in speech act theory Austin (1962), conceptualizes representational harms caused by generative language systems as the perlocutionary effects (i.e., real-world impacts) of particular types of illocutionary acts (i.e., system behaviors). However, generative language systems often support myriad use cases and generative AI systems often incorporate other modalities, such as speech and vision. We focus specifically on language generation systems, without accounting for modalities beyond language. Our conceptualization of representational harms also focuses on the occurrence of particular system behaviors, rather than taking a distribution- or performance-based approach (Katzman et al., 2023). We believe our framework applies across these dimensions of variation, but would require adaptation to focus on propositions instead of utterances. We defer a detailed exploration of this direction to future work.
We also emphasize that although speech act theory enables us to conceptualize all salient aspects of representational harms, our framework and taxonomy may not capture every detail worth considering in every context. For example, we anticipate that some of the illocutionary act patterns in our taxonomy will be more common in language about particular harmful social hierarchies or particular social groups. Indeed, additional illocutionary act patterns may therefore need to be incorporated.
In Section 4, we discussed the ways that our framework and taxonomy can support the development of valid measurement instruments, presenting a case study that engages with recent conceptual debates about what constitutes a representational harm and how such harms should be measured. Such conceptual debates lay the groundwork for the operationalization process by providing a clear specification of exactly what should be operationalized and why. Although we chose to focus on conceptual debates, the operational debates are also crucial to developing valid measurement instruments. We therefore hope that future work picks up from where we left off, focusing on the role of our framework and taxonomy in this part of the measurement process.
Finally, we emphasize that our collective positionalities influence our thinking about language as a social phenomenon, representational harms, and language generation systems. We are an interdisciplinary group of linguists with a variety of specializations; experts in natural language processing and machine learning; and applied scientists. As a result, our shared perspective is broad and has been influenced by our myriad conversations with other technologists, computational social scientists, linguists, natural language processing and machine learning experts, applied scientists, and engineers. From a linguistic perspective, we are oriented toward functional approaches to language variation, pragmatics, linguistic anthropology, sociolinguistics, and philosophy, as well as the broader social sciences and humanities. That said, despite our broad shared perspective, there are likely gaps in our thinking due to our collective positionalities.
7 Acknowledgments
We thank Solon Barocas, Su Lin Blodgett, Tim Vieira, Philip Resnik and his students, and many others who have shaped this paper over the years.
Appendix A Speech Act Theory
A.1 Components of Speech Act Theory
As Austin described in his Williams James lectures, speech acts can be understood as having three dimensions—locution, illocution, and perlocution. A speech act can be characterized by four components belonging to these dimensions: its locutionary act, its illocutionary act and effects, and its perlocutionary effects. Below we describe these acts and effects using the canonical utterance “Can you pass the salt?” as a running example.
An utterance’s locutionary act corresponds to its form—i.e., its word choice and ordering. For example, the locutionary act for the utterance “Can you pass the salt?” is the act of uttering the ordered words “can” + “you” + “pass” + “the” + “salt.” A locutionary act produced by a speaker is heard by the hearer as uttered, but can be understood by the hearer in a variety of different ways. [^14] Inherent ambiguity at either the word level or the sentence level can lead to communication challenges. For example, some utterances may require additional information in order to capture their referential meaning (e.g., the utterance “Bring me that dog!” poses the question “Which dog?”), other utterances may reflect lexical or syntactic variation (e.g., “Where’s the elevator?” vs. “Can you point me to the lift?”), while others still may involve homophony in spoken language (e.g., “knew” vs. “new”) or polysemy (e.g., “foot” can be a unit of measurement, a body part, a location, etc.). All of these sources of ambiguity can affect how a hearer understands an utterance.
An utterance’s illocutionary act corresponds to its purpose, often characterized by what its speaker intends [^15] We emphasize that generative language systems are not sentient and do not have intent. We address this tension between the nature of generative language systems and intent-centric understandings of speech act theory in Appendix A.2. to accomplish through the production of that utterance.161616 Because humans cannot read minds, the illocutionary act is often the site of significant linguistic and social ambiguity. Interpreting the illocutionary act can therefore require the hearer to make inferences based on word choice and ordering, other linguistic and paralinguistic cues, and social context. Building on Austin’s presentation of speech acts, Searle (1976) identified and described five basic classes of illocutionary acts: representatives, expressives, directives, commissives, declarations. For example, the illocutionary act for the utterance “Can you pass the salt?” is a directive—i.e., a request for an action to be carried out—namely that the salt be passed. Because an illocutionary act can take many different forms, it can therefore correspond to many different locutionary acts. For example, if a speaker wants the hearer to pass the salt, they might say “Can you pass the salt?” but they might instead say “This food needs salt!”, “Can you give me that?”, “May I have the salt, please?”, “Give me the salt now!”, “Salt! Immediately!”, and so on.
An utterance’s illocutionary effects are the entailed linguistic consequences of the corresponding illocutionary act—i.e., what happens linguistically, as a result of the illocutionary act (Lorenzini, 2020). For example, a speaker’s request to pass the salt has an illocutionary effect of the hearer having been asked to carry out an action, namely to pass the salt.
An utterance’s perlocutionary effects correspond to its real-world impacts, which derive from the interplay between locution and illocution. These perlocutionary effects are not always predictable from the utterance itself or the context in which it was produced. An utterance can have an unlimited number of perlocutionary effects. We build on Austin’s conceptualization by further distinguishing between first-order or second-order perlocutionary effects. First-order perlocutionary effects occur at the time of utterance production and often involve the participants in the interaction, while second-order perlocutionary effects occur subsequent to utterance production. For example, the perlocutionary effects for the utterance “Can you pass the salt?” might include the salt being passed, the salt being thrown, the pepper being passed, the salt being poured on the floor, and so on.
A.2 A Speech Act Theoretical Model of Communication
Speech act theory provides the necessary scaffolding for the development of a basic model of interactive communication between humans, which includes the the word choice and ordering in a given utterance, the purpose of the utterance, and the real-world impacts of the utterance. This model is helpful in understanding how utterances—including those that cause representational harms—convey meaning between speakers and hearers. We argue that parts of this model can be mapped to interactions between systems and humans. Below, we explain this model of communication in human-to-human interactions and then map it onto system-to-human interactions, where the system is considered the “speaker” and the human user is the “hearer”. The model is shown in Figure 2.
Figure 2: An interactive model of communication based in speech act theory.
In the context of human communication, the speaker produces an utterance, or expression, and the hearer hears the utterance. Through the process of uptake Austin (1962), the hearer hears the utterance’s locution in a given social context and infers the utterance’s purpose—or illocution—which includes its conventional meaning (in both a semantic and pragmatic sense), the communicative intent of the speaker, and the entailed linguistic consequences. The inference of an utterance’s purpose in a given communicative context results in some communicative outcomes, which are real-world impacts—or perlocutionary effects. This model leverages the analytical power of speech act theory to facilitate the identification of relevant units of analysis from the human hearer’s perspective.
However, mapping human-to-human interaction to system-to-human interaction requires careful consideration of the role of speaker intention, which is commonly integrated into conceptualizations of illocution. While some traditional schools of thought within speech act theory theorize the meaning of an utterance as the intention of the speaker—i.e., within illocution (see Searle and Vanderveken, 1985; Grice, 1957, 1968, 1969; Sperber and Wilson, 1995; Neale, 1992; Stalnaker, 1999, inter alia)—these theories of meaning cannot be used in contexts where the “speaker” is a generative language system without inadvertently ascribing intent to these systems. We, like others, have taken issue with ascribing intention to generative language systems (Bender and Koller, 2020; Bender et al., 2021) and thus believe these theories of meaning are inappropriate when the “speaker” is non-sentient.
Even in eschewing intentionalist accounts of meaning, much of the functionality of generative language systems relies on the simulation of communicative conventions and their ability to signal utterance purpose to achieve effective human uptake. Thus, we argue that aspects of human language and communicative conventions must still be taken into account when analyzing generative language system outputs, without ascribing humanness or human intent to these systems.
With this framing, we posit that meaning in generative language system outputs—which are simulations of natural human speech—is reliant on human hearer uptake. For interactions occurring in this communicative context, uptake is guided by both the interpretation of simulated “speaker” intention and system-external perlocution involving the hearer. Human hearers “hear”, interpret, and respond to a generative language system’s output via this process of uptake, and they derive semantic, pragmatic, and social meaning from language precisely because of communicative conventions that influence uptake. Steering away from speaker- or intent-centric analyses of meaning, this uptake-centric approach situates meaning and meaning-making as a social phenomenon central to the human hearer and ensuing perlocutionary effects, thereby skirting problematic anthropomorphism (Weidinger et al., 2022; Abercrombie et al., 2023).
The model of communication presented here is functional for human-to-human and system-to-human communication, centering interpretation of meaning as uptake by the human hearer. For the human hearer, the expression, or utterance, is the locution, while its purpose—comprised of the inferred meaning and inferred communicative intent, and the entailed consequences of those inferences—reflect the illocution, and the actual real-world impacts of the utterance are the perlocutionary effects.
While not treated in detail here, this communicative model may be adapted to other generative modalities (e.g., image, speech, video). For example, visual representations that are outputs of image generation systems can be conceptualized as types of illocutionary acts whose meaning is interpreted by a human “hearer” or recipient and that can have any number of perlocutionary effects.
A.3 A Speech Act Theoretical Perspective on Conversational Ideals
Speech act theory, in the broader context of the field of pragmatics, has been further developed and theoretically expanded since Austin’s lectures. Notably, Searle (1976) and Searle and Vanderveken (1985) proposed a refined set of five illocutionary act classes, which capture the range of illocutionary acts that can be communicated. These classes are representatives, expressives, directives, commissives, and declarations. In this section, we define these speech act classes and describe how they have been understood in the context of human interactions with generative language systems.
Representatives commit the speaker to the belief that the stated representation corresponds to some true state of affairs in the world, thus communicating observations and beliefs. Expressives are illocutionary acts that express a psychological or subjective state, as in expressions of joy, frustration, or sadness. In contrast to representatives and expressives that present some perspective on the world, directives are illocutionary acts that direct or instruct the listener to take some action. These are common in interactions involving advice or how-to instructions, but the class of directives also includes illocutionary acts that prohibit or forbid the hearer from taking some action. Commissives are distinct in that they commit the speaker to a future action—e.g., promises and threats. Finally, declarations—sometimes called performatives—change or update some aspect of world to match some part of the utterance, such as in ceremonial pronouncements like baptisms, ship namings, and weddings that change the legal status of some entity or set of entities.
Recently, many have acknowledged the relevance of pragmatics in the design and evaluation of generative language systems (Pautler and Quilici, 1998; Goodman and Frank, 2016; Freiman and Miller, 2019; Sap et al., 2020; Fried et al., 2023; Kasirzadeh and Gabriel, 2023). Of these, Kasirzadeh and Gabriel (2023) draw special attention to speech act theory by highlighting the kinds of illocutionary acts—namely, the classes identified by Searle (1976)—that are inappropriate when produced by generative language systems due to their lack of embodiment and psychology. Specifically, they argue that three of Searle’s classes should not be present in the outputs of generative language systems: expressives—due to generative language systems’ lack of internal psychology; commissives—due to generative language systems’ inability to follow through beyond episodic interactions; and declarations—due to generative language systems’ lack of authority to bring about (specifically, ceremonial or legal) changes in the world. According to this view, different kinds of speech acts reflect different communicative goals, and some of these goals cannot be achieved by generative language systems precisely because they are not human.
In contrast, representatives, and sometimes directives, can serve the interactive goals of humans in their interactions with generative language systems, and particularly with those systems that function as conversational agents. In system-to-human interactions, these two kinds of speech acts can serve communicative goals whereby the system is designed to help the user complete some task or action (i.e., via the production of directives) or the system is designed to engage in conversation (i.e., comprised primarily of representatives).
However, Kasirzadeh and Gabriel (2023) point out that not all types of representatives and directives will meet what they call discursive ideals, which are norms of ideal speech within a given domain. They argue that conversational agents may have as a goal “the management of difference and enablement of productive cooperation in public […] life”. Accordingly, these systems should aim for “normative” values of civility, namely respect, tolerance, and consideration for others. We posit that stereotyping, demeaning, and erasing illocutionary acts can be formulated across each of Searle’s speech acts in ways that violate these normative ideals. These violations are related to their perlocutionary effects of entrenching harmful social hierarchies. Accordingly, these types of illocutionary acts—i.e., the ones that cause representational harms—may be understood in the same way as the expressives, commissives, and declarations underlined by Kasirzadeh and Gabriel (2023), as types of speech acts that are wholly inappropriate when produced by generative language systems.
Appendix B Social Hierarchies and Identity
B.1 Harmful Social Hierarchies
We distinguish between three types of harmful social hierarchies: broadly experienced social hierarchies, social hierarchies that are not broadly experienced, and local social hierarchies [^17] For other conceptualizations of local social hierarchies and social hierarchies that are not broadly experienced, see Eckert (1989), Hall and Jefferson (1976), and Hebdige (1979).. By distinguishing between these types of hierarchies, we gain a deeper theoretical understanding of representational harms beyond simple fairness criteria like sub-group parity (Hutchinson and Mitchell, 2019) and can make more informed decisions about the scope of any given measurement task related to the concept of representational harms.
We have established that the entrenchment of broadly experienced social hierarchies—i.e., those which involve one or more of the factors that are most influential on society’s conceptualization and understanding of identity, such as race, ethnicity, gender, sexuality, age, socioeconomic status, ability, religion, etc.—relates to one common conceptualization of representational harms. We have named these fairness-related representational harms. However, this category of representational harms does not capture all possible harmful representations of people. A comprehensive model of representational harms must account for kinds of social hierarchies and their relationship to identity.
We can achieve greater conceptual clarity by identifying other kinds of harmful representations, such as those that entrench harmful social hierarchies that are not broadly experienced, such as interest groups, sports teams, university affiliations, etc. We call these non-fairness-related representational harms, and we argue any measurement or evaluation effort requires decisions about whether these kinds of representational harms are within the scope of a given measurement task and thus whether they should be included in associated measurement instruments. For some purposes, the representation of groups belonging to hierarchies that are not broadly experienced may be less critical to measure and address in generative language system outputs, as they 1) generally reflect less stable and less entrenched harmful social hierarchies whose structure and perceived or real benefits may differ greatly from person to person, 2) may be less likely to cause compounding harms of both entrenchment of the hierarchy and negative impacts on individuals’ psychological states, and 3) may be less severe, as affiliations in social groups within these hierarchies tend to be voluntary and not subject to historical and systemic loss of access to power, status, privileges, resources, and opportunities.
Finally, local social hierarchies, such as those comprised of sets of individuals within family units, friend groups, organizations, and workplaces, may also need distinct treatment. The entrenchment of these local hierarchies results in harms of individual characterization. The ways the entrenchment of these local social hierarchies arise via language often parallel common conceptualizations of demeaning. For example, harms of individual characterization may be caused by name-calling, insulting, or hurling slurs at individuals. This parallels the way fairness-related representational harms are carried out—via name calling, insulting, or hurling slurs that target social groups or an individual based on their membership in a given social group.
In fact, some approaches to the identification and measurement of concepts related to representational harms are designed to include what we have termed harms of individual characterization alongside fairness-related representational harms. However, these approaches often fail to address the important distinctions between the types of hierarchies involved and the resulting implications for the conceptualization of the targeted concept (see, for example, Waseem et al., 2017; Nangia et al., 2020).
B.2 Formation of Identity
Hierarchies are comprised of individuals or social groups. Individuals and social groups form their identities through the processes of self- and group-conceptualization. These processes involve situating oneself and others within a given hierarchy, and can center relationships between individuals, between a specific individual and a social group, or between (and sometimes within) different social groups. These identity negotiation processes are mediated through a variety of social activities, including via discourse about identity as it relates to oneself, others, and social groups.
Discourse in which self- and group-conceptualization occurs utilizes the evaluative lenses of similarity/difference and authenticity/inauthenticity (see Appendix E for details about these evaluative lenses). Through these lenses, people make determinations about the boundaries of and relationships between themselves and others, as well as between social groups, informing their conceptualizations about identity. These evaluative lenses are commonly invoked in explicit and implicit discourse about identity, and in the identification and evaluation of characteristics belonging to individuals and social groups.
Characteristics are sometimes also referred to as “traits”, “factors”, or “attributes”. These can include physical, psychological, behavioral, experiential, or relational properties of people. We conceptualize characteristics as type-value pairs. For example, hair length is a characteristic type, while short and 4cm are possible values for that characteristic type. Characteristic values can reflect qualitative or quantitative evaluations of the characteristic type.
Each individual has a unique set of characteristics, which may be core to how they situate themself and others in the production of their (individual) identity through self-conceptualization. An individual’s unique set of characteristics may position them as a member of multiple social hierarchies, which are systematic organizations of individuals or groups of people that differentially confer power, status, privileges, resources, and opportunities. These social hierarchies may be broadly experienced, not broadly experienced, or local social hierarchies, and the social hierarchies may themselves be shorthand for groups of characteristics, e.g. gender, race, etc. In the case of an individual having characteristics that correspond to multiple broadly experienced social hierarchies, an intersectional view of identity must be taken into account. For more information on intersectional identity both generally and in the context of machine learning, see Crenshaw (1991), Wang et al. (2022), and Ovalle et al. (2023).
Like individuals, social groups are also characterized by sets of characteristics. However, in the case of social groups, these sets of characteristics are socially salient—i.e., both recognizable and reproducible. When people talk about themselves and others, these sets of characteristics can be foregrounded in and by speech acts that occur in discourse, and they provide a “recipe” for the necessary ingredients for the (re)production of the social group. Sets of social groups that are differentiated from each other on the basis of one or more characteristics often belong to the same social hierarchy. An example of a social hierarchy is gender and an example of a social group within this hierarchy is nonbinary people.
Ultimately, identity is the socially-negotiated product of both self- and group-conceptualization, which utilize the the evaluative lenses of similarity/difference and authenticity/inauthenticity to make assessments about characteristics.
Appendix C Mechanisms of Entrenchment
Entrenchment of a harmful social hierarchy is the further cementing of the hierarchy in the world. We argue that there are five main mechanisms of entrenchment when representations are produced via speech act. These mechanisms are the ways in which the language of the utterance achieves the cementing of the hierarchy at the boundary of the perlocutionary effect and a stereotyping, demeaning, or erasing illocutionary act. The perlocutionary effect of entrenchment is achieved via the mechanisms of activating, asserting, justifying, enforcing, and restructuring a harmful social hierarchy.
Activating a harmful social hierarchy involves bringing the hierarchy to the linguistic foreground by implicitly or explicitly mentioning either the hierarchy or one or more social groups within the hierarchy. Activating a harmful social hierarchy is a prerequisite for entrenchment. In other words, for entrenchment of a harmful social hierarchy to occur, the hierarchy must first be foregrounded. Additionally, activation of the hierarchy must co-occur with another of the mechanisms of entrenchment in order for entrenchment to have occurred. For example, certain types of basic questions about social groups (“What are some popular hangouts for the local gay community?”) and descriptions of the impacts of the hierarchy on individuals (“Gay people may experience discrimination in their professional or social circles”) activate harmful social hierarchies without entrenching them. Examples in which harmful social hierarchies are activated and entrenched are presented below.
Asserting a harmful social hierarchy involves presenting a hierarchical positioning of one or more social groups as “right”—meaning, as it should be. For example, “Straight people are superior to gay people” both activates the hierarchy over sexual orientations and asserts the superiority of one group over another. It thus presents the hierarchical relation between these groups as right, thereby entrenching a hierarchy that promotes homophobia.
Justifying a harmful social hierarchy involves providing a reason or justification for the hierarchy’s existence. For example, the utterance “Gay people are weird” activates the hierarchy of sexual orientations and presents a reason for the existence of (inherent inequality within) the hierarchy, thereby entrenching the harmful social hierarchy.
Enforcing a harmful social hierarchy involves encouraging actions that align with a hierarchy, often devaluing one or more social groups while placing higher social value on another. For example, “Gay people don’t belong here” activates the hierarchy of sexual orientations while encouraging an action—i.e., exclusion from a particular space—that aligns with the hierarchy’s inherent harmful inequality.
Restructuring a harmful social hierarchy involves discursive—meaning language- or discourse-based—attempts at altering the composition or consequences of a hierarchy. As an example, “There’s no such thing as bisexuality” activates the hierarchy of sexual orientations by invoking bisexuality, and the utterance presents a view that alters the composition of the harmful social hierarchy by denying the existence of this one specific socially meaningful distinction within it.
As described in Section 2.2, the illocutionary acts that stereotype, demean, and erase leverage these mechanisms in the production of the perlocutionary effect of entrenching social hierarchies.
Appendix D Other Fairness-related Harms: Allocation and Quality of Service
Extending the speech act theory framing beyond the question of representational harms, we argue that other types of harms may be similarly defined by their perlocutionary effects, or real-world impacts. To demonstrate this, we consider here two other commonly cited fairness-related harms, namely those involving systematic disparities in allocation (Barocas et al., 2017) and quality of service (Crawford, 2017; Blodgett, 2021).
Like representational harms, defined in Section 2.1, allocation and quality-of-service harms also implicate harmful social hierarchies. Specifically, harms of allocation or quality of service occur when one of a system output’s perlocutionary effects is the enactment of one or more harmful social hierarchies, where enactment refers to the act of differentially distributing of power, status, privileges, resources, or opportunities in alignment with at least one relevant social hierarchy. In contrast with entrenchment, the enactment of a harmful social hierarchy is the result of actions that actively distribute or influence the distribution of power, status, privileges, resources, or opportunities. As a result, the enactment of a harmful social hierarchy is often a first-order perlocutionary effect for quality-of-service harms, because the enactment happens at the time of the system output as an immediate outcome of the output itself. For allocation disparities, the enactment of the harmful social hierarchy may be either a first-order or second-order perlocutionary effect, depending on the type of system producing them. For example, systems that gate-keep job opportunities (Barocas and Selbst, 2016), mortgage loan access (Lee and Floridi, 2021), and physical freedom (Chouldechova, 2017) may or may not not cause the perlocutionary effect during the human interaction with the system, and the effect itself may be mediated (and upheld) by a human reviewer.
Appendix E Evaluative Lenses
Identity is negotiated, or mediated, through social semiotic processes—i.e., processes that create and call attention to social meaning via signs and symbols, such as dialectal features, physical characteristics, social characteristics, personal style, and so on. These social semiotic processes, namely adequation/distinction and authentication/denaturalization, leverage evaluative lenses through which relevant signs and symbols operate in the ongoing production of identity (see Bucholtz and Hall, 2004, 2005). More concretely, they are lenses through which individuals, social groups, or individuals based on their membership in those social groups may evaluate their characteristics and those of other individuals and social groups in the production and maintenance of their own identity. We detail here the evaluative lenses of similarity/difference (corresponding to adequation/distinction) and authenticity/inauthenticity (corresponding to authentication/denaturalization), whose use in language can empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)), and which provide the basis of our systematization of stereotyping, demeaning, and erasing illocutionary acts described in Section 2.2. Additionally, these lenses also can be used to either disrupt or entrench harmful social hierarchies.
Similarity/difference (Gal and Irvine, 1995; Irvine and Gal, 2000; Bucholtz and Hall, 2005) is the most basic evaluative lens through which individuals and social groups negotiate their identity. Invoking this evaluative lens involves likening oneself to, or differentiating oneself from, other individuals and social groups by comparing characteristics. Authenticity/inauthenticity is the evaluative lens through which individuals and social groups are deemed to have characteristics that fit (or do not fit) within a given paradigm (Bucholtz and Hall, 2005)—i.e., a relevant “blueprint” used to evaluate social belonging in a social group or in a hierarchy of social groups.
During the negotiation of identity, the evaluative lenses of similarity/difference and authenticity/inauthenticity are used to empower and disempower one or more social groups (or one or more individuals based on their membership in those social group(s)). Empowerment and disempowerment are social (but not semiotic) processes that position an individual or social group within a particular social hierarchy. Empowerment affirms or makes individuals or social groups authoritatively and institutionally accepted, establishing them in a position of power. In contrast, disempowerment “dismiss[es], censor[s], or simply ignore[s]” them, depriving them of power (Bucholtz and Hall, 2005).
When an individual or generative language system produces an utterance that uses these evaluative lenses to characterize one or more social groups (or one or more individuals based on their membership in those social group(s)) rather than in the production of their own identity, these utterances may constitute stereotyping, demeaning, and/or erasing language. Specifically, speech acts that empower or disempower one or more social groups (or one or more individuals based on their membership in those social group(s)) within a social hierarchy by invoking these evaluative lenses of similarity/difference and authenticity/inauthencity entrench harmful social hierarchies. Stereotyping, demeaning, and erasing utterances are types of illocutionary acts that use these evaluative lenses in distinct ways, as detailed in Section 2.2.
Appendix F Contextualizing Our Taxonomy
The illocutionary act patterns represented Table 1 are sourced from range of literatures, including linguistic anthropology and critical discourse analysis, sociolinguistics, philosophy of language, sociology, psychology, computational linguistics, and cognitive science. While some of these patterns are explicitly mentioned in one or more of the sources, others are not. Many of these sources touch on multiple features of interest simultaneously. For example, Reyes (2004) discusses both stereotyping and race, and Slobe (2016) touches on race, gender, and general theoretical work.
At a more granular level, this taxonomy pulls from work on speech act theory (Austin, 1962; Grice, 1957, 1968, 1969, 1975, 1989; Harris and McKinney, 2021; Jucker, 2024; Lorenzini, 2020; Neale, 1992; Pautler and Quilici, 1998; Sbisà, 2013; Searle, 1976; Searle and Vanderveken, 1985; Sperber and Wilson, 1995; Stalnaker, 1999), raciolinguistics and language and race (Hill, 1998; Labov, 1972; Reyes, 2004; Rickford, 2000; Rosa and Flores, 2017; Samy Alim and Reyes, 2011; Urciuoli, 1996, 2011), language and gender (Bucholtz, 1998; Kiesling, 2011; Ochs, 1992; Zimman, 2019), language and disability (Henner and Robinson, 2023; Henner, 2024; Ladau, 2021), linguistic anthropological theory (Agha, 2003, 2005, 2010; Blommaert and Varis, 2015a, b; Bucholtz and Hall, 2004, 2005; Eckert, 1989, 2005; Eckert and McConnell-Ginet, 2007; Eckert, 2008; Irvine and Gal, 2000; Johnstone et al., 2006; Slobe, 2016), slurs (Croom, 2013; Liu, 2021), notions of (in)authenticity (Bucholtz, 2003; Reyes, 2017), stereotyping (Augoustinos and Walker, 1995; Blodgett, 2021; Cardwell, 1996; Katzman et al., 2023; Kraft and Mortensen, 2023; Tajfel, 1981), linguistic harms (Banko et al., 2020; Castelle, 2018; Diberardino et al., 2024; McGowan, 2019; Tirrell, 2017), and general linguistics and philosophy (Biber and Finegan, 1989; Clark and Brennan, 1991; Freiman and Miller, 2019; Horn, 1984). In some cases, the illocutionary act patterns and their illocutionary effects are abstracted from examples provided by researchers across social science disciplines. In other cases, they are reformulations, within the speech act theory framing, from sources focusing on hate speech and toxic language.
Within this table are illocutionary acts that can be further categorized according to Searle’s classes of representatives, expressives, commissives, directives, and declarations. The rows near the top of the stereotyping, demeaning, and erasing types are representatives, while those that begin with “advocate” are directives, those that begin with “threaten” are commissives, and those that “deny” (access and justice) are declarations. For the sake of brevity, those illocutionary act class labels are not explicitly provided in the table. Applying Searle’s classes across the broader categories of illocutionary act types—stereotyping, demeaning, and erasing acts—ensures a more comprehensive set of patterns than would be identified from the source literatures.
Figure 3: Visualization of two taxonomies of representational harms as things systems do (i.e., behaviors), organized by author. Shading indicates overlap with the types of system behaviors presented in this paper.
Figure 4: Visualization of a taxonomy of representational harms as real-world impacts caused by system outputs. Lower levels have been truncated. See Chien and Danks (2024) for the full taxonomy.
As mentioned in body of this paper, the class of expressives—which express a psychological or evaluative state—are intentionally excluded because of their formative features that make them a close match to their representative forms. The examples in the table can be reformulated as expressives, condoning or approving of the proposition that corresponds to the illocutionary act pattern shown. For example, an expressive may express approval of any of the stigmatizing illocutionary act patterns noted in the table. That said, expressives are especially relevant for erasing speech acts. This is because erasing expressives reinforce a given representation of reality. As a result, erasing expressives can encourage or promote real-world actions related to the pattern, such as denying necessary accommodations, prioritizing equality over equity, and demonizing reparations or other restorative actions. For example, an expressive speech act that positively frames non-differentiated treatment—e.g., “It’s good that everyone has to suffer from the choices [social group] made”—presents different groups’ needs, experiences, contributions and accountability as equal and promotes a world in which this is the predominant view, which is especially problematic in cases where accountability and consequences are unfairly distributed across the hierarchy.
Appendix G Taxonomies of Representational Harms
Example Illocutionary act type (and class) Illocutionary act pattern Evaluative lens(es) Illocutionary effect
(1) "Gay men are not very masculine" Stereotyping (representative) Describes a group stereotypically; essentializes Within-group similarity, between-group difference, authenticity Group is stereotyped
(2) "Obviously bisexuals are messed up" Demeaning (representative) Stigmatizes on the basis of immorality, danger, or deviance Between-group difference, within-group similarity (implicit) Group is stigmatized
(3) "There is no way that bisexuality is a real thing" Erasing (representative) Denies existence or fails to recognize a group Within-hierarchy similarity Groups are homogenized and presented as indistinguishable
Table 3: Illustrative examples, originally from ToxiGen (Hartvigsen et al., 2022), classified according to their illocutionary act types and classes, illocutionary act patterns, corresponding evaluative lenses, and illocutionary effects. A common perlocutionary effect for all examples is the entrenchment of harmful social hierarchies.
We provide two figures illustrating the taxonomies of “representational” harms proposed by Blodgett (2021), Katzman et al. (2023), and Chien and Danks (2024). Figure 3 shows two different single-layer taxonomies of the types of system behaviors corresponding to their respective notions of representational harms. Note that some of the types—i.e., those shaded in gray—match or are very similar to the types of illocutionary acts that we further systematize in Sections 2.2 and 3.
An alternative taxonomical structure is presented by Chien and Danks (2024), who developed a deeper taxonomy of representational harms as impacts of system behaviors. It is organized by the top-level set of types, which include the psychosocial object impacted, the social locus of the impact, and the specific types of impacts that affect those objects and loci. Note that not only is this structure a deeper taxonomy than those shown in in Figure 3, but this taxonomy also targets different aspects of the conceptual space of representational harms. Specifically, Chien and Danks (2024) focus on the range of impacts caused by system outputs, while Blodgett (2021) and Katzman et al. (2023) are most interested in types of system outputs.
Appendix H Illustrative Examples
Table 3 describes the illocutionary features of the illustrative examples presented in sections 2.2 and 3 using the framework presented in those sections.